OpenStudy (anonymous):

please help me c: ! Any line with an undefined slope is parallel to the......... x-axis, y-axis, or x=y ?

OpenStudy (e.mccormick):

Slope is change in y over change in x. Undefined is a way of saying divide by 0. So undefined slope is change in y over 0! So, which do you think that would be?

OpenStudy (anonymous):

x=y ?

OpenStudy (e.mccormick):

\[\frac{y_1-y_2}{0}\]Means that \(x_1-x_2=0\)

OpenStudy (e.mccormick):

Not x=y, but what would cause the bottom numbers to cancel each other?

OpenStudy (primeralph):

@_AriannaMoniquee Got it?

OpenStudy (anonymous):

y-axis right ?

OpenStudy (e.mccormick):

Yah. Cause \(x_1-x_2=0\) means the xs are all the same!

OpenStudy (e.mccormick):

The parllel to x-axis is when all the ys are the same, or 0 slope. So that is how you can keep track of these. No change in y and it is parallel to x, but no change in x and it is parallel to y.

OpenStudy (anonymous):

ohh! i understand now , thank u c:

OpenStudy (anonymous):

yes @primeralph

OpenStudy (e.mccormick):

np. Have fun!